Let {vn} be a sequence in R2
, say vn = (xn, yn). Give R2
the || ||∞
norm. Show that limn→∞ vn → v if and only if limn→∞ xn = x and
limn→∞ yn = y where v = (x, y).
1
Expert's answer
2020-06-22T17:40:08-0400

Given that "\{ v_n\}" be a sequence in "\R^2" .

Let "\{ v_n\}=(x_n,y_n)" and "v=(x,y)" .


Let "lim_{n\to \infty} v_n=v"

Claim : "x_n\rightarrow x \ and \ y_n\rightarrow y \ as \ n\rightarrow \infty"

i,e "lim_{n\to \infty} x_n=x \ and \ lim_{n\to \infty} y_n=y" .

As "lim_{n\to \infty} v_n=v" ,so for a given "\epsilon >0" there exist a "K\in \N"

such that "||v_n-v||_{\infty}<\epsilon \ \forall n \geq K"

"\implies ||x_n-x,y_n-y||_{\infty}<\epsilon \ \forall n \geq K"

"\implies max\{ |x_n-x|,|y_n-y|\}<\epsilon" "\forall n \geq K"

"\implies |x_n-x|<\epsilon \ and \ |y_n-y|<\epsilon \" "\forall n\geq K"

Hence "lim_{ n\to \infty} x_n=x \ and \ lim_{n\to \infty} y_n=y."

Conversely , assume that "lim_{n\to \infty}x_n=x \ and \ lim_{ n\to \infty}y_n=y"

Claim: "lim_{n\to \infty} v_n=v" .

Therefore , for any "\epsilon >0 , \exist \ N_1 , N_2" such that

"| x_n-x|<\epsilon" for all "n\geq N_1"

and "|y_n-y|<\epsilon" for all "n\geq N_2" .

Now , "||v_n-v||_{\infty}=||x_n-x,y_n-y||_{\infty}"

"=Max \{ |x_n-x|,|y_n-y|\} <\epsilon" "\forall n\geq N"

Where "N=Max\{ N_1,N_2\}" .

Hence , "lim_{n\to \infty}v_n=v."